OpenStudy (anonymous):

Find the area of a triangle with a base length of 3 units and a height of 4 units. a 3 units b 6 units c 7 units d 8 units

OpenStudy (ajanijones):

Hint to find the Area of a triangle its, Height*Base/2 So 4*3 = x Then x/2

OpenStudy (ajanijones):

So its 4*3 = 12 then 12/2 = 6 So 6 is your final answer.